OpenStudy (anonymous):

Solve the following at interval [0,2pi) 7+5sinx-cos^2 x=0

OpenStudy (anonymous):

start with \[7+5\sin(x)-(1-\sin^2(x))=0\] then solve the resulting quadratic equation in \(\sin(x)\)